THE ARCHIVE

INSTITUTE FOR MATHEMATICAL SCIENCES UNIVERSITY OF COPENHAGEN

φ (\frac{∑ a_{ν}x_{ν}}{∑ a_{ν}} ) ≤ \frac{∑ a_{ν}φ(x_{ν})}{∑ a_{ν}}

May 3rd, 1998

Rules governing the use of archival material at the Archive

Permission to examine

Permission to examine archival material is granted to scholars upon completion of the registration form. Such permissions are granted subject to whatever restrictions may have been placed on the material by their donors, depositors or the archive, which may then limit the access to certain collections or part of collections. As a general rule archival material more than 30 years old will be accessible. In certain cases, scholars may be required to use microfilm, printed copies or an electronic version of the material when such copies are available. Remark that permission to examine archival material does not automatically give permission to quote from the material (see the section “Publication and quotation” below).

Rules for handling the archival material

All archival material must be handled with great care and according to the following rules:

• Food, drink, and smoking are prohibited in the archive room or a possible other room where archival material is studied.

• To prevent reordering or mixing of the material only one box must be opened and only the content of one folder must be examined at a time.

• All archival material have to be returned in exactly the same condition as it was received, which means that it is not allowed to:

1. use archival material as writing pad,

2. put opened books or booklets on top of each other,

3. rearrange the order of materials in folders or boxes (even though the order does not seem logical),

4. to lean on or write on the archival material,

5. to fold the material anew,

6. to moisten your fingers during the work.

Photocopying

Usually visitors of the archive will get permission to photocopy archival material themselves by an oral inquiry under the condition that the photocopies are not to be used for any purpose other than private study, scholarship, or research. However, the archive may refuse to give such a permission if, in its judgement, this would damage the material or involve violation of copyright law or restriction of the specific material. During photocopying one has to take the condition of the material into consideration.

Photocopies of specific archival material may be ordered. Such orders are usually done by a secretary or by students, and all expenses (salary, paper etc) have to be paid by the scholar.

Publication and quotation

Permission to examine archival material is not an authorization to publish them. To the extent that it may do so legally, the archive ordinarily will grant publication right for scholarly purposes upon written request by qualified applicants. If the archive is not able to grant the right for publication or quotation it will help you to find where such right can be obtained. Dissertations and theses are considered published works. Researchers who plan to publish their work are urged to make inquiry concerning possible restrictions before beginning their research.

In giving permission to publish a manuscript, or portions of a manuscript, the Archive does not surrender its own right thereafter to publish the manuscript or grant permission to others to publish it.

If permission to publish is granted, the location of the cited material shall be indicated in the published work. Oftentimes citations from archival material in published books and articles include an incomplete reference, making it difficult for the researcher and the archivists to find the document to which the author refers. For the benefit of other researchers, we would like to remind you to include the following information in your citations to documents from the archive:

1. Title or description of documents, including author, recipient or audience, and date. For reports, include formal title, preferable in quotation marks or italic.

2. Folder number. Where material is unprocessed or no folder number appears, use a possible title of the folder.

3. Box number.

4. Number and name of series and record groups. This only applies to large collections.

5. Name of the collection.

6. The Archive, Institute for Mathematical Sciences, University of Copenhagen, Denmark. Optionally also the address: Universitetsparken 5, DK-2100 Copenhagen. The full identification only needs to be given after the first quotation; afterwards an abbreviated form may be used.

Examples of correct references:

• Letter from G. H. Hardy to H. Bohr March 25th, 1912, folder 8, box 2, Harald Bohr Papers; The Archive, Institute for Mathematical Sciences, University of Copenhagen, Denmark (below abbreviated: AIMS).

• Carl C. Hansen: Hvorledes har vort Kendskab til Funktionen ζ(s) udviklet sig siden Riemanns Tid? [How has our knowledge of the function ζ(s) developed since the time of Riemann?]. Opgave til Magisterkonferens i Matematik, Maj-Juni 1899. Box 2, Carl C. Hansen papers, AIMS.